ON THE NONEXISTENCE OF SYMMETRIC BLOCK DESIGN (81,16,3)

WITH INVOLUTORY HOMOLOGY by Vladimir Ćepulić and Ljubo Marangunić, Zagreb, Croatia

Abstract. We show that (81,16,3)-block design admitting involutory homology does not exists. Mathematics subject classification: 05B05 Keywords: Symmetric block design, automorphism group, involutory homology, orbital structure.